# Spin-up and spin distribution of stellar black holes grown by gas accretion in proto-stellar clusters

### 19 Mar 2026 v1

@@
11institutetext: Dipartimento di Fisica “G. Occhialini”, Universitá degli Studi di Milano-Bicocca, Piazza della Scienza 3, 20126 Milano, Italy 22institutetext: Istituto Nazionale di Fisica Nucleare (INFN), Sezione di Milano-Bicocca, Piazza della Scienza 3, 20126 Milano, Italy
@@
Proto-stellar clusters, likely progenitors of globular clusters (GCs), are extremely compact with typical masses $\sim 10^{6}\,{\rm M}_{\odot}$ and sizes $\sim 1\,{\rm pc}$, as revealed recently by James Webb Space Telescope (JWST) observations at $z\sim 10$. Sufficiently high compactness can provide a time window for early-formed stellar black holes (BHs) to accrete primordial gas. We develop a semi-analytic model to follow BH spin-up and determine the final spin distribution of stellar BHs which grow in mass via gas accretion within compact gaseous proto-stellar clusters. The velocity shear within a BH’s sphere of influence induces the formation of an accretion disk which is repeatedly disrupted by stochastic perturbations to the BH motion. We assume low initial BH spins $a_{*,{\rm ini}}=0.01$, consistent with stellar-evolution models with efficient angular-momentum transport, and we restrict initial BH masses to values below the upper BH mass gap, $m_{\rm BH,ini}<55\,{\rm M}_{\odot}$. Our analysis shows a strong BH spin-mass correlation, obtained within $\sim 10\,{\rm Myr}$ when gas is depleted. Low-spin BHs, $a_{*}\leq 0.3$, are predominantly low-mass, $m_{\rm BH}\lesssim 25\,{\rm M}_{\odot}$, in contrast to high-spin black holes, $a_{*}\geq 0.7$, which are predominantly high-mass, $m_{\rm BH}\gtrsim 65\,{\rm M}_{\odot}$. Notably, there exist also low-spin, high-mass outliers with $\sim 1$ mass-gap BH per cluster expected to have $a_{*}\sim 0.1$. The general trend, however, expressed by the median spin as a function of final BH mass is well fit by a high-spin saturating exponential with transition mass $\sim 50\,{\rm M}_{\odot}$. For $m_{\rm BH}\geq 100\,{\rm M}_{\odot}$ the median spin is $\bar{a}_{*}\sim 0.90$ with the central $68\%$ of the distribution spanning $a_{*}\sim 0.70-0.96$, in striking agreement with the estimated spins of the BH components of the gravitational-wave signal GW231123. These spin values persist up to the highest masses generated by our mechanism, $m_{\rm BH}\sim 10^{3}\,{\rm M}_{\odot}$.
@@ 1 Introduction
In Roupas (2025) (hereafter Paper I) we proposed a channel for the BH mass growth via gas accretion in proto-stellar clusters to values within the upper BH mass gap, induced by the physics of a pair-instability supernova (PISN) (Belczynski et al., 2016; Spera and Mapelli, 2017; Woosley and Heger, 2021), $60\,{\rm M}_{\odot}\lesssim m_{\rm BH}\lesssim 130\,{\rm M}_{\odot}$, and up to the intermediate-massive-black-hole (IMBH) regime of $m_{\rm BH}\sim 10^{3}\,{\rm M}_{\odot}$. Gas-rich, compact, low-metallicity, massive stellar clusters provide ideal conditions for the mass growth of stellar BHs via accretion. Mass segregation, which drives the massive stellar progenitors of BHs toward the dense center of the cluster (Spitzer, 1987), is accelerated in the presence of ambient gas remaining from the first stellar formation event, at timescales $\lesssim 1\,{\rm Myr}$ (Leigh et al., 2014). Moreover, in a lower-metallicity environment stellar winds get weaker, allowing for longer gas-depletion timescales in the cluster, despite PISN feedback, for sufficiently compact clusters (Paper I).
@@
Recent observations by JWST revealed five proto-stellar cluster candidates in the Cosmic Gems arc galaxy at high redshift, $z=10.2$, (Adamo et al., 2024), which present low metallicity and are remarkably compact. Their typical masses are $\sim 10^{6}\,{\rm M}_{\odot}$ and half-light radii $\sim 1\,{\rm pc}$. Such proto-stellar clusters are plausibly the progenitors of globular clusters (Kruijssen, 2015; Renzini, 2017; Elmegreen, 2018). These JWST observations provide observationally motivated cluster parameters for our models.
@@
The semi-analytic framework of stellar-BH growth in such environments, introduced in Paper I, incorporates gas depletion by stellar feedback, distinct BH formation times depending on stellar progenitors and formation channel, dynamical friction, stochastic stellar encounters, turbulence, and gas accretion. Here we extend this model to include the BH spin evolution. This calculation is essential for identifying our proposed channel regarding the astrophysical origin of gravitational-wave (GW) signals, among different stellar-BH mass growth channels. These include the repeated BH mergers channel (Miller and Hamilton, 2002; O’Leary et al., 2006; Gerosa and Berti, 2017), and gas accretion in AGN disks (Levin, 2007; McKernan et al., 2012; Yi et al., 2018; Yang et al., 2020), as well as in the cores of proto-galactic haloes (Safarzadeh and Haiman, 2020; Bartos and Haiman, 2025).
@@
Identifying the origin of heavy BHs $\sim 10^{2}-10^{3}\,{\rm M}_{\odot}$ is a timely, active area of research (Gerosa and Berti, 2017; Roupas and Kazanas, 2019; Baibhav et al., 2020; Renzo et al., 2020; Kimball et al., 2020; Safarzadeh and Haiman, 2020; Mapelli et al., 2021; Rizzuto et al., 2022; Arca Sedda et al., 2023; Torniamenti et al., 2024; Fujii et al., 2024; Purohit et al., 2024; Baumgarte and Shapiro, 2025; Lee et al., 2025; Bartos and Haiman, 2025; Kıroğlu et al., 2025; Gottlieb et al., 2025; Liu et al., 2025). The GW experiments, LIGO-Virgo-KAGRA Collaborations, have revealed a significant population of BHs within the upper BH mass gap (Abbott and others, 2023). The most recent such GW signal, GW231123 (Abac et al., 2025), involves components which are both heavy, $103_{-52}^{+20}\,{\rm M}_{\odot}$, $137_{-17}^{+22}\,{\rm M}_{\odot}$, and have high spins, $0.9_{-0.19}^{+0.10}$, $0.8_{-0.51}^{+0.20}$. These values are difficult to explain within the repeated-merger scenario, which predicts a narrow peak at $\sim 0.7$ (Fishbach et al., 2017). We show here that these mass and spin values naturally emerge via accretion in gas-rich proto-stellar clusters.
@@
The BH motion is repeatedly deflected by stochastic gravitational perturbations –two-body encounters and turbulent gas motions– that reorient the BH trajectory. Angular-momentum inflow is induced by the velocity shear within the BH’s sphere of influence, which can produce an accretion disk. These successive episodes lead to an accretion regime, where a broad range of spin values is possible, $\sim 0-0.97$, with a strong spin-mass correlation as we demonstrate below.
@@
In the next section we review the cluster and BH motion model of Paper I. In Section 3 we describe the BH spin evolution and accretion disk models. Section 4 presents our results, and we conclude in Section 5.
@@ 2 Cluster physical model
We adopt the same semi-analytic model we introduced in Paper I for the proto-stellar cluster and the BH motion. We summarize briefly the main elements in this section and refer the reader to Paper I for further details.
@@ 2.1 Gas depletion timescale
rather than on the cluster mass alone. A simple theoretical understanding follows from noting that the binding energy scales as $U\propto M^{2}/R\sim C\cdot M$, and therefore increases linearly with the total mass for a fixed initial compactness, while the total stellar feedback energy scales as $E_{\rm feed}\propto M$ regardless of compactness. Consequently, compactness –not mass– primarily determines the relative energy balance between feedback and gravitational binding. The star formation efficiency is also crucial in setting the depletion timescale, as it regulates both the available gas reservoir and the total energy released by stellar feedback. A third key parameter is metallicity, which influences stellar wind luminosities and stellar lifetimes, thereby the onset of SN feedback. In this work we focus on low-metallicity clusters ($Z\sim 0.01Z_{\odot}$) motivated by high redshift JWST observations of the Cosmic Gems arc galaxy.
In Fig. 1, we display the depletion timescale $\tau$ as a function of compactness for several values of $\varepsilon$. This calculation is based on our detailed model presented in Paper I, where corresponding figures for solar and sub-solar metallicities are also provided. In our model $\tau$ is defined as the time when the cumulative energy from stellar winds and SN equals the binding energy. We took into account the stellar lifetimes dependence on stellar masses, extrapolating the tabulated data, referring to low-metallicity stars, of Groh et al. (2019). Stellar-wind and SN energy injection is modelled based on a carefull consideration of observations and numerical data (Vink et al., 2001; Kudritzki, 2002; Puls et al., 2008; Krause et al., 2013; Renzo et al., 2017; Sander and Vink, 2020). We observe in Fig. 1 a characteristic stall at
@@
which marks the onset of pair-instability SNe (PISNe) for the adopted parameters. The range of compactness values for the observed proto-stellar clusters in the Cosmic Gems arc galaxy –the same range used in our simulations– corresponds to this stall timescale.
@@ 2.2 Black hole motion
In order to simulate the motion of BHs, we implemented the same model of Paper I briefly summarized as follows. We assumed background gas and stellar Plummer density profiles taking into account the expansion of the cluster due to the mass loss, the higher extension of the gas component relative to the stellar one, the gas temperature profile and the gas cooling over time. We adopted an exponential law of gas depletion (Baumgardt and Kroupa, 2007),
@@
since the stellar feedback processes, responsible for gas depletion, generate energy proportional to the cluster mass. We terminate each run once $99\%$ of the initial gas mass has been depleted. In practice, the BH mass and spin evolution stall even earlier, when $\sim 95\%$ of the gas has been removed.
At each simulation run, we generated a BH population from a Salpeter BH initial mass function (Perna et al., 2019) sampling positions and velocities from the Brownian probability density of Roupas (2021). Each BH was created at a different time depending on the zero-age main sequence (ZAMS) progenitor mass. We classified the BHs on several types depending on the ZAMS progenitors mass assumed to drive the related formation channel: core-collapse-SN (CCSN), direct-collapse of low or high ZAMS mass, and pulsational-pair-instability-SN (PPISN) of low or high ZAMS mass. We applied a recoil kick at birth to the BH remnants of CCSN which has been injecting $\sim 70\%$ of such BHs out of the cluster.
Figure 2: The contour of the median stochastic timestep $\tilde{\Delta t}_{\rm stoch}$ with respect to the median BH position $\tilde{r}_{\rm BH}$ and the final BH mass $m_{\rm BH,f}$ for an indicative run of our typical cluster $M_{\star}=10^{6}\,{\rm M}_{\odot}$, $r_{c,\star}=1\,{\rm pc}$, $\varepsilon=0.35$.
We evolved all BHs in the evolving background gas and stellar profiles integrating the BHs equations of motion with a standard drift-kick scheme. The deterministic forces – gravity and dynamical friction from the background– were implemented using an adaptive Runge-Kutta method, while stochastic velocity increments due to both stellar perturbations and gas turbulencewe were applied at adaptive timesteps, $\Delta t_{\rm stoch}$, as described in detail in our Paper I. The local gravitational dynamical timescale $\tau_{\rm grav}(r_{\bullet})=\sqrt{r_{\bullet}^{3}/(GM_{\rm enc}(r_{\bullet}))}\sim 1/\sqrt{G\left\langle\rho_{\rm tot}(r_{\bullet})\right\rangle}$ is the typical orbital timescale of the local perturber population. It is therefore a natural choice for setting the reference scale of the timestep $\Delta t_{\rm stoch}$ for applying stochastic kicks. We further impose consistency bounds discussed in detail in Appendix A. In Figure 2 we plot the contour of the median $\tilde{\Delta t}_{\rm stoch}$ with respect to the median BH position and the final BH mass for an indicative run of our typical cluster. It is $\tilde{\Delta t}_{\rm stoch}=\mathscr{O}(10^{-2}\,{\rm Myr})$ and can become as low as $\sim 0.007\,{\rm Myr}$ for more heavy BHs, $m_{\bullet,{\rm f}}\gtrsim 50\,{\rm M}_{\odot}$, which move on average deeper in the cluster core.
We adopt an isotropic hot-type accretion, whose typical representative is Bondi-Hoyle accretion (Merritt, 2013) given the dense, hot, turbulent gaseous environment. The spherically symmetric accretion rate is (Frank et al., 2002)
@@
The appropriate value of accretion radius $R_{\rm acc}$ depends on the relative amount of gas and the radial position of the accretor (Bonnell et al., 2001). If the gas dominates the cluster potential at a certain BH position, $r_{\bullet}$, the accretion rate is determined by a tidal-lobe accretion radius (Paczyński, 1971)
@@
We assume the accretion radius to be the smaller of the two at any instant in time (Bonnell et al., 2001),
@@
adopting the most conservative assumption. Furthermore, we adopted a maximum threshold for accretion rate
$\dot{m}_{\rm acc}(t)=\min\left\{\dot{m}_{\rm feed}(t),10\dot{m}_{{\rm Edd},0}\right\}$ (8)
where $\dot{m}_{{\rm Edd},0}=4\pi Gm_{\rm BH}m_{p}/\eta_{0}\sigma_{T}c$ is the reference Eddington rate value for the radiative efficiency $\eta_{0}=0.057$. This is based on the results of Inayoshi et al. (2016) who have shown that for $1<\dot{m}_{\rm feed}/\dot{m}_{\rm Edd,0}<100$ the mean accretion rate cannot exceed $\left\langle\dot{m}_{\rm acc}\right\rangle\lesssim 10\dot{m}_{\rm Edd,0}$. We adopted this constant cap as a simple, conservative assumption, since Inayoshi et al. (2016) report that there occur episodic bursts of accretion rate comparable to $\dot{m}_{\rm feed}$ violating the general cap for the mean rate. In practice, in the physical conditions we consider here and for $m_{\rm BH}\lesssim 10^{3}{\rm M}_{\odot}$ we have $0<\dot{m}_{\rm feed}(t)/\dot{m}_{\rm Edd,0}<10$, satisfying automatically the cap.

### 29 Apr 2026 v2

@@
11institutetext: Dipartimento di Fisica “G. Occhialini”, Universitá degli Studi di Milano-Bicocca, Piazza della Scienza 3, 20126 Milano, Italy 11email: zacharias.roupas@unimib.it 22institutetext: Istituto Nazionale di Fisica Nucleare (INFN), Sezione di Milano-Bicocca, Piazza della Scienza 3, 20126 Milano, Italy
@@
Proto-stellar clusters, likely progenitors of globular clusters, are extremely compact with typical masses of $\sim 10^{6}\,{\rm M}_{\odot}$ and sizes of $\sim 1\,{\rm pc}$, as has recently been revealed by James Webb Space Telescope observations at $z\sim 10$. Sufficiently high compactness can provide a time window for early-formed stellar black holes (BHs) to accrete primordial gas. We developed a semi-analytic model to follow BH spin-up and determine the final spin distribution of stellar BHs that grow in mass via gas accretion within compact gaseous proto-stellar clusters. The velocity shear within a BH’s sphere of influence induces the formation of an accretion disk that is repeatedly disrupted by stochastic perturbations to the BH motion. We assumed low initial BH spins of $a_{*,{\rm ini}}=0.01$, consistent with stellar-evolution models with efficient angular-momentum transport, and we restricted initial BH masses to values below the upper BH mass gap, $m_{\rm BH,ini}<55\,{\rm M}_{\odot}$. Our analysis shows a strong BH spin-mass correlation, obtained within $\sim 10\,{\rm Myr}$ when gas is depleted. Low-spin BHs, $a_{*}\leq 0.3$, are predominantly low-mass, $m_{\rm BH}\lesssim 25\,{\rm M}_{\odot}$, in contrast to high-spin BHs, $a_{*}\geq 0.7$, which are predominantly high-mass, $m_{\rm BH}\gtrsim 65\,{\rm M}_{\odot}$. Notably, there exist also low-spin, high-mass outliers with $\sim 1$ mass-gap BH per cluster expected to have $a_{*}\sim 0.1$. The general trend, however, expressed by the median spin as a function of final BH mass, is well fit by a high-spin saturating exponential with a transition mass of $\sim 50\,{\rm M}_{\odot}$. For $m_{\rm BH}\geq 100\,{\rm M}_{\odot}$ the median spin is $\bar{a}_{*}\sim 0.90$, with the central $68\%$ of the distribution spanning $a_{*}\sim 0.70-0.96$, in striking agreement with the estimated spins of the BH components of the gravitational-wave signal GW231123. These spin values persist up to the highest masses generated by our mechanism, $m_{\rm BH}\sim 10^{3}\,{\rm M}_{\odot}$.
@@ 1 Introduction
In Roupas (2025) (hereafter Paper I) we proposed a channel for the black hole (BH) mass growth via gas accretion in proto-stellar clusters to values within the upper BH mass gap, induced by the physics of a pair-instability supernova (PISN) (Belczynski et al., 2016; Spera and Mapelli, 2017; Woosley and Heger, 2021), $60\,{\rm M}_{\odot}\lesssim m_{\rm BH}\lesssim 130\,{\rm M}_{\odot}$, and up to the intermediate-massive-black-hole (IMBH) regime of $m_{\rm BH}\sim 10^{3}\,{\rm M}_{\odot}$. Gas-rich, compact, low-metallicity, massive stellar clusters provide ideal conditions for the mass growth of stellar BHs via accretion. Mass segregation, which drives the massive stellar progenitors of BHs toward the dense center of the cluster (Spitzer, 1987), is accelerated in the presence of ambient gas remaining from the first stellar formation event, at timescales of $\lesssim 1\,{\rm Myr}$ (Leigh et al., 2014). Moreover, in a lower-metallicity environment stellar winds get weaker, allowing for longer gas-depletion timescales in the cluster, despite PISN feedback, for sufficiently compact clusters (Paper I).
@@
Recent observations by the James Webb Space Telescope (JWST) revealed five proto-stellar cluster candidates in the Cosmic Gems arc galaxy at high redshift, $z=10.2$, (Adamo et al., 2024), which present low metallicity and are remarkably compact. Their typical masses are $\sim 10^{6}\,{\rm M}_{\odot}$ and half-light radii $\sim 1\,{\rm pc}$. Such proto-stellar clusters are plausibly the progenitors of globular clusters (GCs) (Kruijssen, 2015; Renzini, 2017; Elmegreen, 2018). These JWST observations provide observationally motivated cluster parameters for our models.
@@
The semi-analytic framework of stellar-BH growth in such environments, introduced in Paper I, incorporates gas depletion by stellar feedback, distinct BH formation times depending on stellar progenitors and formation channel, dynamical friction, stochastic stellar encounters, turbulence, and gas accretion. Here we extend this model to include the BH spin evolution. This calculation is essential for identifying our proposed channel regarding the astrophysical origin of gravitational-wave (GW) signals, among different stellar-BH mass growth channels. These include the repeated BH merger channel (Miller and Hamilton, 2002; O’Leary et al., 2006; Gerosa and Berti, 2017), and gas accretion in active galactic nucleus (AGN) disks (Levin, 2007; McKernan et al., 2012; Yi et al., 2018; Yang et al., 2020) as well as in the cores of proto-galactic haloes (Safarzadeh and Haiman, 2020; Bartos and Haiman, 2025).
@@
Identifying the origin of heavy BHs $\sim 10^{2}-10^{3}\,{\rm M}_{\odot}$ is a timely, active area of research (Gerosa and Berti, 2017; Roupas and Kazanas, 2019; Baibhav et al., 2020; Renzo et al., 2020; Kimball et al., 2020; Safarzadeh and Haiman, 2020; Mapelli et al., 2021; Rizzuto et al., 2022; Arca Sedda et al., 2023; Torniamenti et al., 2024; Fujii et al., 2024; Purohit et al., 2024; Baumgarte and Shapiro, 2025; Lee et al., 2025; Bartos and Haiman, 2025; Kıroğlu et al., 2025; Gottlieb et al., 2025; Liu et al., 2025). The GW experiments, LIGO-Virgo-KAGRA Collaborations, have revealed a significant population of BHs within the upper BH mass gap (Abbott and others, 2023). The most recent such GW signal, GW231123 (Abac et al., 2025), involves components that are heavy, $103_{-52}^{+20}\,{\rm M}_{\odot}$, $137_{-17}^{+22}\,{\rm M}_{\odot}$, and that have high spins, $0.9_{-0.19}^{+0.10}$, $0.8_{-0.51}^{+0.20}$. These values are difficult to explain within the repeated-merger scenario, which predicts a narrow peak at $\sim 0.7$ (Fishbach et al., 2017). We show here that these mass and spin values naturally emerge via accretion in gas-rich proto-stellar clusters.
@@
The BH motion is repeatedly deflected by stochastic gravitational perturbations – two-body encounters and turbulent gas motions – that reorient the BH trajectory. Angular-momentum inflow is induced by the velocity shear within the BH’s sphere of influence, which can produce an accretion disk. These successive episodes lead to an accretion regime, where a broad range of spin values, $\sim 0-0.97$, is possible, with a strong spin-mass correlation, as we demonstrate below.
@@
In the next section we review the cluster and BH motion model of Paper I. In Sect. 3 we describe the BH spin evolution and accretion disk models. Section 4 presents our results, and we conclude in Sect. 5.
@@ 2 Cluster physical model
We adopted the same semi-analytic model as we introduced in Paper I for the proto-stellar cluster and the BH motion. We briefly summarize the main elements in this section and refer the reader to Paper I for further details.
@@ 2.1 Gas depletion timescale
rather than on the cluster mass alone. A simple theoretical understanding follows from noting that the binding energy scales as $U\propto M^{2}/R\sim C\cdot M$, and therefore increases linearly with the total mass for a fixed initial compactness, while the total stellar feedback energy scales as $E_{\rm feed}\propto M$ regardless of compactness. Consequently, compactness – not mass – primarily determines the relative energy balance between feedback and gravitational binding. The star formation efficiency is also crucial in setting the depletion timescale, as it regulates both the available gas reservoir and the total energy released by stellar feedback. A third key parameter is metallicity, which influences stellar wind luminosities and stellar lifetimes, and thereby the onset of SN feedback. In this work we focus on low-metallicity clusters ($Z\sim 0.01Z_{\odot}$) motivated by high-redshift JWST observations of the Cosmic Gems arc galaxy.
In Fig. 1, we display the depletion timescale, $\tau$, as a function of compactness for several values of $\varepsilon$. This calculation is based on our detailed model presented in Paper I, where corresponding figures for solar and sub-solar metallicities are also provided. In our model $\tau$ is defined as the time when the cumulative energy from stellar winds and SN equals the binding energy. We took into account the stellar lifetimes dependence on stellar masses, extrapolating the tabulated data, referring to low-metallicity stars, of Groh et al. (2019). Stellar-wind and SN energy injection was modeled based on a careful consideration of observations and numerical data (Vink et al., 2001; Kudritzki, 2002; Puls et al., 2008; Krause et al., 2013; Renzo et al., 2017; Sander and Vink, 2020). We observe in Fig. 1 a characteristic stall at
@@
that marks the onset of PISNe for the adopted parameters. The range of compactness values for the observed proto-stellar clusters in the Cosmic Gems arc galaxy – the same range as is used in our simulations – corresponds to this stall timescale.
@@ 2.2 Black hole motion
In order to simulate the motion of BHs, we implemented the same model as in Paper I, briefly summarized as follows. We assumed background gas and stellar Plummer density profiles, taking into account the expansion of the cluster due to the mass loss, the higher extension of the gas component relative to the stellar one, the gas temperature profile, and the gas cooling over time. We adopted an exponential law of gas depletion (Baumgardt and Kroupa, 2007),
@@
since the stellar feedback processes, responsible for gas depletion, generate energy proportional to the cluster mass. We terminated each run once $99\%$ of the initial gas mass had been depleted. In practice, the BH mass and spin evolution stalled even earlier, when $\sim 95\%$ of the gas had been removed.
At each simulation run, we generated a BH population from a Salpeter BH initial mass function (Perna et al., 2019) sampling positions and velocities from the Brownian probability density of Roupas (2021). Each BH was created at a different time depending on the zero-age main sequence (ZAMS) progenitor mass. We classified the BHs into several types depending on the ZAMS progenitor mass assumed to drive the related formation channel: a core-collapse-SN (CCSN), direct collapse of low or high ZAMS mass, and pulsational-pair-instability-SN (PPISN) of low or high ZAMS mass. We applied a recoil kick at birth to the BH remnants of CCSN that injected $\sim 70\%$ of such BHs out of the cluster.
Figure 2: Contour of the median stochastic timestep, $\tilde{\Delta t}_{\rm stoch}$, with respect to the median BH position, $\tilde{r}_{\rm BH}$, and the final BH mass, $m_{\rm BH,f}$, for an indicative run of our typical cluster, $M_{\star}=10^{6}\,{\rm M}_{\odot}$, $r_{c,\star}=1\,{\rm pc}$, $\varepsilon=0.35$.
We evolved all BHs in the evolving background gas and stellar profiles, integrating the BHs equations of motion with a standard drift-kick scheme. The deterministic forces – gravity and dynamical friction from the background – were implemented using an adaptive Runge-Kutta method, while stochastic velocity increments due to both stellar perturbations and gas turbulencewe were applied at adaptive timesteps, $\Delta t_{\rm stoch}$, as is described in detail in Paper I. The local gravitational dynamical timescale, $\tau_{\rm grav}(r_{\bullet})=\sqrt{r_{\bullet}^{3}/(GM_{\rm enc}(r_{\bullet}))}\sim 1/\sqrt{G\left\langle\rho_{\rm tot}(r_{\bullet})\right\rangle}$, is the typical orbital timescale of the local perturber population. It is therefore a natural choice for setting the reference scale of the timestep, $\Delta t_{\rm stoch}$, for applying stochastic kicks. We further imposed consistency bounds, which are discussed in detail in Appendix A. In Fig. 2 we plot the contour of the median $\tilde{\Delta t}_{\rm stoch}$ with respect to the median BH position and the final BH mass for an indicative run of our typical cluster. It is $\tilde{\Delta t}_{\rm stoch}=\mathscr{O}(10^{-2}\,{\rm Myr})$ and can become as low as $\sim 0.007\,{\rm Myr}$ for more heavy BHs, $m_{\bullet,{\rm f}}\gtrsim 50\,{\rm M}_{\odot}$, which move on average deeper in the cluster core.
We adopted an isotropic hot-type accretion, whose typical representative is Bondi-Hoyle accretion (Merritt, 2013) given the dense, hot, turbulent gaseous environment. The spherically symmetric accretion rate is (Frank et al., 2002)
@@
The appropriate value of the accretion radius, $R_{\rm acc}$, depends on the relative amount of gas and the radial position of the accretor (Bonnell et al., 2001). If the gas dominates the cluster potential at a certain BH position, $r_{\bullet}$, the accretion rate is determined by a tidal-lobe accretion radius (Paczyński, 1971),
@@
We assumed the accretion radius to be the smaller of the two at any instant in time (Bonnell et al., 2001),
@@
adopting the most conservative assumption. Furthermore, we adopted a maximum threshold for the accretion rate of
$\dot{m}_{\rm acc}(t)=\min\left\{\dot{m}_{\rm feed}(t),10\dot{m}_{{\rm Edd},0}\right\},$ (8)
where $\dot{m}_{{\rm Edd},0}=4\pi Gm_{\rm BH}m_{p}/\eta_{0}\sigma_{T}c$ is the reference Eddington rate value for the radiative efficiency, $\eta_{0}=0.057$. This is based on the results of Inayoshi et al. (2016), who showed that for $1<\dot{m}_{\rm feed}/\dot{m}_{\rm Edd,0}<100$ the mean accretion rate cannot exceed $\left\langle\dot{m}_{\rm acc}\right\rangle\lesssim 10\dot{m}_{\rm Edd,0}$. We adopted this constant cap as a simple, conservative assumption, since Inayoshi et al. (2016) reported that there occur episodic bursts of accretion rate comparable to $\dot{m}_{\rm feed}$ that violate the general cap for the mean rate. In practice, in the physical conditions we consider here and for $m_{\rm BH}\lesssim 10^{3}{\rm M}_{\odot}$ we have $0<\dot{m}_{\rm feed}(t)/\dot{m}_{\rm Edd,0}<10$, automatically satisfying the cap.